# What Expressivity Theory Misses: Message Passing Complexity for GNNs

### v1

@@ Abstract
Expressivity theory, characterizing which graphs a GNN can distinguish, has become the predominant framework for analyzing GNNs, with new models striving for higher expressivity. However, we argue that this focus is misguided: First, higher expressivity is not necessary for most real-world tasks as these tasks rarely require expressivity beyond the basic WL test. Second, expressivity theory’s binary characterization and idealized assumptions fail to reflect GNNs’ practical capabilities. To overcome these limitations, we propose Message Passing Complexity (MPC): a continuous measure that quantifies the difficulty for a GNN architecture to solve a given task through message passing. MPC captures practical limitations like over-squashing while preserving the theoretical impossibility results from expressivity theory, effectively narrowing the gap between theory and practice. Through extensive validation on fundamental GNN tasks, we show that MPC’s theoretical predictions correlate with empirical performance, successfully explaining architectural successes and failures. Thereby, MPC advances beyond expressivity theory to provide a more powerful and nuanced framework for understanding and improving GNN architectures.
@@ 1 Introduction
From weather forecasting to drug design, Graph Neural Networks (GNNs) have shown remarkable success across diverse applications. However, the seminal works by Morris et al. [2019] and Xu et al. [2019] revealed a key limitation of standard Message Passing Neural Networks (MPNNs): Their ability to distinguish non-isomorphic graphs can be bounded by the Weisfeiler-Leman (WL) graph isomorphism test Weisfeiler and Leman [1968].
@@
In response, significant research effort has focused on developing more expressive architectures that surpass the WL test Jegelka [2022]. The underlying hypothesis is that increased expressivity translates to better empirical performance, with improved benchmark results often attributed to higher expressivity Wollschläger et al. [2024]. Isomorphism-based (Iso) expressivity theory, which characterizes the sets of graphs an architecture can distinguish through the WL test and its extensions, has thus become the predominant framework for analyzing MPNNs.
@@
We argue that this focus on iso expressivity is misguided. While it provides valuable impossibility results, we identify two key limitations that prevent it from explaining real-world MPNN performance. First, higher expressivity is often not necessary for real-world tasks: almost all graphs in standard benchmarks are already distinguishable by the basic WL test, making it unclear why higher expressivity would improve performance Zopf [2022]. Second, iso expressivity theory fails to capture practical model capabilities. It assumes unrealistic conditions like lossless information propagation over unbounded layers, ignoring practical limitations like over-squashing Alon and Yahav [2021]. Moreover, its binary view (can vs. cannot distinguish) offers no insight into the relative difficulty of learning specific real-world tasks.
@@
To address these limitations, we propose shifting from binary expressivity tests to a continuous complexity measure, MPC, that quantifies the message-passing complexity of arbitrary tasks for a given architecture. MPC builds upon a novel probabilistic WL test. It captures practical limitations of MPNNs, such as under-reaching Barceló et al. [2020] and over-squashing Alon and Yahav [2021], which are known to hinder empirical performance, while preserving impossibility results from iso expressivity theory—effectively narrowing the gap between theory and practice.
@@
Through extensive validation, we show that trends in MPC complexity align with empirical performance across a range of fundamental graph tasks. Notably, success is determined not by iso expressivity but by architectural choices that minimize complexity for specific tasks. For instance, a simple GCN with virtual node outperforms strictly more expressive, higher-order models at long-range tasks, as MPC correctly predicts. By providing a quantitative measure of architectural capabilities for specific tasks, MPC both reveals current model limitations and offers clear optimization targets for future architectural innovations, shifting focus from maximizing expressivity to minimizing task-specific complexity.
@@ 2 Background
Here, $\smash{{\texttt{msg}}^{l}_{0}}$, $\smash{{\texttt{msg}}^{l}_{1}}$, and $\smash{{\texttt{upd}}^{l}}$ can be arbitrary (learned) functions, often MLPs. Typical choices for the aggregation function agg are mean or sum. We differentiate between a model architecture $\mathcal{M}$} that only specifies which nodes exchange messages (abstracting away from the choice of msg, upd, and agg), and a (learned) model instantiation ${M}\in{\mathcal{M}}$ with fully specified msg, upd and agg functions. Standard MPNNs ${\mathcal{M}}_{S}$} perform message passing directly on the input graph G{G}. In contrast, more recent architectures propagate messages on a transformed message passing (MP) graph $\smash{{\tilde{{G}}}}=t({G})$ Velickovic [2022]. They introduce modifications like additional virtual nodes Gilmer et al. [2017], rewired edges Topping et al. [2022], or higher-order graphs Wollschläger et al. [2024]. For simplicity, we mainly focus on standard MPNNs ${\mathcal{M}}_{S}$} in the main text, and defer a general framework encompassing architectures $\mathcal{M}$} with arbitrary MP graphs $\smash{{\tilde{{G}}}}$ to §˜B.1.
@@ 3 Limitations of Expressivity Theory
MPNN architectures differ in their theoretical capacity to solve graph tasks, such as detecting specific substructures Zhang et al. [2024]. Most prior expressivity theory captures differences by considering an architecture’s ability to distinguish non-isomorphic graphs relative to a reference isomorphism test $\alpha$ Giusti et al. [2023], Bodnar et al. [2021a], Wollschläger et al. [2024], Morris et al. [2021]. Formally:
@@
Standard MPNNs ${\mathcal{M}}_{S}$} are at most as expressive as the WL test Morris et al. [2019], Xu et al. [2019]. In response, recent GNN research has focused on developing architectures that surpass the WL test in expressivity. Their strong performance on benchmarks such as ZINC Bodnar et al. [2021a], Bouritsas et al. [2023], Feng et al. [2023], Wollschläger et al. [2024] is often motivated by and attributed to their higher iso expressivity Wollschläger et al. [2024], Bouritsas et al. [2023], Feng et al. [2023]. This builds on the premise that theoretical iso expressivity correlates with empirical performance. Consequently, this line of expressivity research rests on two fundamental assumptions, both of which we show to be problematic (Fig.˜1):
@@
Figure 1: Limitations of Iso expressivity: 1) Expressivity does not accurately describe practical capabilities of GNNs, and 2) limited expressivity does not restrict performance in real-world tasks. MPC addresses all limitations narrowing the gap between theory and practice.
@@
We first critique Assumption 1, showing that iso expressivity theory relies on idealized assumptions and provides incomplete insights about practical MPNN capabilities. Positive expressivity results only guarantee the existence of a model instance within the architecture that is more expressive than $\alpha$. These proofs typically rely on a maximally expressive model $M^{*}$ with injective upd functions and an unbounded number of layers Bodnar et al. [2021a]: conditions under which every node can access the complete information from all other nodes. This idealized setup ignores fundamental limitations of real-world MPNNs. In practice, message passing is lossy: information is often bottlenecked (over-squashing Alon and Yahav [2021]), blurred (over-smoothing Li et al. [2018]), or simply out of reach due to shallow depth (under-reaching Barceló et al. [2020]). As a result, real-world MPNNs often fail to propagate information even between nearby nodes, let alone replicate the idealized behavior assumed in expressivity theory. Second, expressivity theory provides strictly binary results Morris et al. [2024]: can vs. cannot distinguish, with no indication of difficulty. This binary lens cannot account for the wide variation in empirical performance between architectures of equal expressivity on the same task. For example, adding a virtual node does not increase iso expressivity Wollschläger et al. [2024] but often leads to performance gains on long-range tasks Cai et al. [2023]. Additionally, between theoretically solvable tasks, iso expressivity theory provides no insight into their relative difficulty. Some tasks might be trivially learnable in practice while others may be practically impossible to learn with finite data and training time. In summary, the assumptions behind positive expressivity results diverge sharply from the realities of real-world MPNNs, and the binary view can offer only limited insights for practice. Our empirical results in §˜5 reinforce this, showing that even highly expressive MPNN architectures struggle with elementary tasks like maintaining initial node features. This highlights a gap between positive iso expressivity results and practical performance.
@@
We now challenge Assumption 2, asking whether negative expressivity results reflect meaningful limitations in practice. Iso expressivity theory considers only the hardest possible task: distinguishing all non-isomorphic graphs. However, many real-world applications, such as social network analysis or relational learning, often require only aggregating (local) information. In these cases, the inability to distinguish certain non-isomorphic graphs does not constitute a meaningful limitation in practice. By focusing exclusively on this worst-case task, expressivity theory overlooks the specific requirements of practical problems. Second, expressivity theory makes global statements over the set of graphs $\mathcal{G}$$^{*}$, demonstrating a lack of expressivity through single, carefully constructed counterexamples Wollschläger et al. [2024], Giusti et al. [2023], Bodnar et al. [2021a]. However, these counterexamples are rare: the probability of encountering WL equivalent graphs in random graphs approaches 0 Babai et al. [1980], and real-world graphs often carry rich node features that further break WL equivalences. Hence, the mere existence of theoretical counterexamples provides limited insight into a model’s practical performance on large, diverse datasets. Lastly, iso expressivity theory is graph-family agnostic. Its results are derived for the set of all possible graphs $\smash{{\mathcal{G}}^{*}}$, whereas real-world applications typically involve restricted graph families. For example, molecular datasets primarily contain planar graphs with bounded degrees. Tasks infeasible over all graphs may become solvable for standard MPNNs within these restricted graph families Dimitrov et al. [2023], Pellizzoni et al. [2025]. Moreover, as shown in Tab.˜1, the basic WL test already distinguishes almost all graph pairs in popular benchmarks across different domains Zopf [2022]. Therefore, iso expressivity theory cannot explain why expressivity beyond the WL test would benefit these real-world tasks or account for the performance differences between architectures.
@@ 4 Message Passing Complexity
To move beyond the limitations of expressivity theory, we propose our continuous, task-specific message passing complexity (MPC). Unlike iso expressivity theory, which only asks whether an MPNN can distinguish certain graphs in theory, MPC quantifies how difficult it is for messages to propagate through a graph to solve a given task. By accounting for the inherently lossy information propagation of real-world MPNNs, MPC can explain practical performance trends that expressivity theory misses. While many complexity measures exist in machine learning and theoretical computer science, they fail to address this unique challenge of MPNNs: propagating information effectively across graph structures. MPC specifically isolates this from other sources of difficulty. For instance, learning high-degree polynomials is difficult but well-understood and not specific to graph learning. For clarity, we focus on node-level tasks $f_{v}$ and standard MPNNs ${\mathcal{M}}_{S}$} in this section, with generalizations to broader architectures and tasks provided in §˜B.1.
@@
Intuitively, $\alpha(x)\vDash_{X}\beta(x)$ means that $\alpha(x)$ provides sufficient information to uniquely determine $\beta(x)$. Using this, we can define a first complexity measure dependent on a specific task $f_{v}$ and individual graph G{G}. The complexity should be maximal if a task $f_{v}$ is infeasible for all $M\in{{\mathcal{M}}_{S}}$, i.e., the WL coloring ${\text{WL}}^{L}_{v}({G})$ of $v$ in the graph G{G} contains insufficient information to deduce $f_{v}({G})$:
@@
$\textup{WLC}_{{\mathcal{M}}_{S}}(f_{v},{G})=\begin{cases}\infty&\text{if }{\text{WL}}^{L}_{v}({G})\nvDash_{{\mathcal{G}}_{\mathcal{V}}}f_{v}({G})\\ 0&\text{else}\end{cases}$
@@ 4.1 Weisfeiler & Leman Go Lossy
To extend our complexity measure beyond the binary characterization of the standard WL test, we need to account for the varying difficulty arising from the lossy message passing observed in practice. This difficulty inherently relies on the graph topology, i.e., the difficulty of propagating a message from one node $u$ to another node $v$ depends on 1) the number of $L$-length walks connecting $v$ and $u$ and 2) the degrees of the nodes on the walk. This can be formalized as the random walk probability ${{\bm{I}}}_{vu}^{L}$ from $v$ to $u$ with edge probabilities $\bm{I}$}} and has been connected to the amount of gradient information node $v$ receives from $u$ Xu et al. [2018] and to the oversquashing phenomenon Giovanni et al. [2023]. Correspondingly, ${{\bm{I}}}_{vu}^{L}$ serves as a measure of difficulty for the simple task of propagating a message from $u$ to $v$.
How can we adapt this measure of difficulty to arbitrary tasks? For this, we propose lossyWL, a probabilistic variant of the WL test that models the possibility of message loss.
Figure 2: Update step of lossyWL for node $v$. Every message $m^{l}_{a\to v}$ (green) survives independently with probability ${{\bm{I}}}_{va}$. lossyWL models the lossy message propagation of real-world MPNNs.
This allows us to quantify difficulty even for WL-distinguishable real-world graphs. Specifically, a message $m_{u\to v}^{l}$ sent from a node $u$ to its neighbor $\smash{v\in{N}_{G}(u)}$ survives with probability $\smash{{{\bm{I}}}_{vu}}$ (Fig.˜2). As a result, the color $\smash{\textup{lossyWL}_{v}^{l}}$ of a node $v$ at iteration $l$ becomes a random variable rather than a deterministic value.
@@
Let $\alpha:X\to Y$ be a probabilistic function, whereas $\beta:X\to Z$ is a deterministic function. Then we can represent $\alpha$ as a deterministic function $\alpha_{s}(x)$ where $s$ is a seed drawn uniformly at random from a set $S$. For a fixed $x\in X$, define:
@@
MPC is based on the probability that under lossy message passing on G{G}, the output after $L$ layers contains sufficient information to deduce $f_{v}$. Intuitively, a high MPC value indicates that the task $f_{v}$ requires 1) combining information from many nodes through 2) messages of low probability, such as messages through bottlenecks. Conversely, a low MPC value means the task depends on information that remains reliably accessible even under lossy message passing conditions. Note that MPC implicitly also depends on $\mathcal{G}$}, which we omit for simplicity. While we focus here on standard MPNNs ${\mathcal{M}}_{S}$}, MPC generalizes naturally to a wide range of architectures $\mathcal{M}$} by performing lossyWL on modified message passing graphs (see §˜B.1). All in all, MPC effectively addresses the discussed limitations of iso expressivity theory, thereby narrowing the gap between theory and practice.

### v2

@@ Abstract
Expressivity theory, characterizing which graphs a GNN can distinguish, has become the predominant framework for analyzing GNNs, with new models striving for higher expressivity. However, we argue that this focus is misguided: First, higher expressivity is not necessary for most real-world tasks as these tasks rarely require expressivity beyond the basic WL test. Second, expressivity theory’s binary characterization and idealized assumptions fail to reflect GNNs’ practical capabilities. To overcome these limitations, we propose Message Passing Complexity (MPC): a continuous measure that quantifies the difficulty for a GNN architecture to solve a given task through message passing. MPC captures practical limitations like over-squashing while preserving theoretical impossibility results from expressivity theory, effectively narrowing the gap between theory and practice. Through extensive validation on fundamental GNN tasks, we show that MPC’s theoretical predictions correlate with empirical performance, successfully explaining architectural successes and failures. Thereby, MPC advances beyond expressivity theory to provide a more powerful and nuanced framework for understanding and improving GNN architectures.
@@ 1 Introduction
From weather forecasting to drug design, Graph Neural Networks (GNNs) have shown remarkable success across diverse applications. However, the seminal works by Morris et al. [36] and Xu et al. [48] revealed a key limitation of standard Message Passing Neural Networks (MPNNs): Their ability to distinguish non-isomorphic graphs can be bounded by the Weisfeiler-Leman (WL) graph isomorphism test [45].
@@
In response, significant research effort has focused on developing more expressive architectures that surpass the WL test [27]. The underlying hypothesis is that increased expressivity translates to better empirical performance, with improved benchmark results often attributed to higher expressivity [46]. Isomorphism-based (Iso) expressivity theory, which characterizes the sets of graphs an architecture can distinguish through the WL test and its extensions, has thus become the predominant framework for analyzing MPNNs.
@@
We argue that this focus on iso expressivity is misguided. While it provides valuable impossibility results, we identify two key limitations that prevent it from explaining real-world MPNN performance. First, higher expressivity is often not necessary for real-world tasks: almost all graphs in standard benchmarks are already distinguishable by the basic WL test, making it unclear why higher expressivity would improve performance [54]. Second, iso expressivity theory fails to capture practical model capabilities. It assumes unrealistic conditions like lossless information propagation over unbounded layers, ignoring practical limitations like over-squashing [1]. Moreover, its binary view (can vs. cannot distinguish) offers no insight into the relative difficulty of learning specific real-world tasks.
@@
To address these limitations, we propose shifting from binary expressivity tests to a continuous complexity measure, MPC, that quantifies the message-passing complexity of arbitrary tasks for a given architecture. MPC builds upon a novel probabilistic WL test. It captures practical limitations of MPNNs, such as under-reaching [3] and over-squashing [1], which are known to hinder empirical performance, while preserving impossibility results from iso expressivity theory—effectively narrowing the gap between theory and practice.
@@
Through extensive validation, we show that trends in MPC complexity align with empirical performance across a range of fundamental graph tasks. [^1] Find our implementation at https://www.cs.cit.tum.de/daml/message-passing-complexity/ Notably, success is determined not by iso expressivity but by architectural choices that minimize complexity for specific tasks. For instance, a simple GCN with virtual node outperforms strictly more expressive, higher-order models at long-range tasks, as MPC correctly predicts. By providing a quantitative measure of architectural capabilities for specific tasks, MPC both reveals current model limitations and offers clear optimization targets for future architectural innovations, shifting focus from maximizing expressivity to minimizing task-specific complexity.
@@ 2 Background
Here, $\smash{{\texttt{msg}}^{l}_{0}}$, $\smash{{\texttt{msg}}^{l}_{1}}$, and $\smash{{\texttt{upd}}^{l}}$ can be arbitrary (learned) functions, often MLPs. Typical choices for the aggregation function agg are mean or sum. We differentiate between a model architecture $\mathcal{M}$} that only specifies which nodes exchange messages (abstracting away from the choice of msg, upd, and agg), and a (learned) model instantiation ${M}\in{\mathcal{M}}$ with fully specified msg, upd and agg functions. Standard MPNNs ${\mathcal{M}}_{S}$} perform message passing directly on the input graph G{G}. In contrast, more recent architectures propagate messages on a transformed message passing (MP) graph $\smash{{\tilde{{G}}}}=t({G})$ [42]. They introduce modifications like additional virtual nodes [18], rewired edges [41], or higher-order graphs [46]. For simplicity, we mainly focus on standard MPNNs ${\mathcal{M}}_{S}$} in the main text, and defer a general framework encompassing architectures $\mathcal{M}$} with arbitrary MP graphs $\smash{{\tilde{{G}}}}$ to §˜B.1.
@@ 3 Limitations of Expressivity Theory
MPNN architectures differ in their theoretical capacity to solve graph tasks, such as detecting specific substructures [53]. Most prior expressivity theory captures differences by considering an architecture’s ability to distinguish non-isomorphic graphs relative to a reference isomorphism test $\alpha$ [22, 5, 46, 37]. Formally:
@@
Standard MPNNs ${\mathcal{M}}_{S}$} are at most as expressive as the WL test [36, 48]. In response, recent GNN research has focused on developing architectures that surpass the WL test in expressivity. Their strong performance on benchmarks such as ZINC [5, 8, 16, 46] is often motivated by and attributed to their higher iso expressivity [46, 8, 16]. This builds on the premise that theoretical iso expressivity correlates with empirical performance. Consequently, this line of expressivity research rests on two fundamental assumptions, both of which we show to be problematic (Fig.˜1):
@@
Figure 1: Limitations of iso expressivity vs. benefits of MPC. Top: Iso expressivity gives an idealized, binary view that misses GNNs’ practical capabilities. Bottom: Limited expressivity rarely restricts real-world performance as it focuses on worst-case graphs and tasks.
@@
We first critique Assumption 1, showing that iso expressivity theory relies on idealized assumptions and provides incomplete insights about practical MPNN capabilities. Positive expressivity results only guarantee the existence of a model instance within the architecture that is more expressive than $\alpha$. These proofs typically rely on a maximally expressive model $M^{*}$ with injective upd functions and an unbounded number of layers [5]: conditions under which every node can access the complete information from all other nodes. This idealized setup ignores fundamental limitations of real-world MPNNs. In practice, message passing is lossy: information is often bottlenecked (over-squashing [1]), blurred (over-smoothing [33]), or simply out of reach due to shallow depth (under-reaching [3]). As a result, real-world MPNNs often fail to propagate information even between nearby nodes, let alone replicate the idealized behavior assumed in expressivity theory. Second, expressivity theory provides strictly binary results [38]: can vs. cannot distinguish, with no indication of difficulty. This binary lens cannot account for the wide variation in empirical performance between architectures of equal expressivity on the same task. For example, adding a virtual node does not increase iso expressivity [46] but often leads to performance gains on long-range tasks [9]. Additionally, between theoretically solvable tasks, iso expressivity theory provides no insight into their relative difficulty. Some tasks might be trivially learnable in practice while others may be practically impossible to learn with finite data and training time. In summary, the assumptions behind positive expressivity results diverge sharply from the realities of real-world MPNNs, and the binary view can offer only limited insights for practice. Our empirical results in §˜5 reinforce this, showing that even highly expressive MPNN architectures struggle with elementary tasks like maintaining initial node features. This highlights a gap between positive iso expressivity results and practical performance.
@@
We now challenge Assumption 2, asking whether negative expressivity results reflect meaningful limitations in practice. Iso expressivity theory considers only the hardest possible task: distinguishing all non-isomorphic graphs. However, many real-world applications, such as social network analysis or relational learning, often require only aggregating (local) information. In these cases, the inability to distinguish certain non-isomorphic graphs does not constitute a meaningful limitation in practice. By focusing exclusively on this worst-case task, expressivity theory overlooks the specific requirements of practical problems. Second, expressivity theory makes global statements over the set of graphs $\mathcal{G}$$^{*}$, demonstrating a lack of expressivity through single, carefully constructed counterexamples [46, 22, 5]. However, these counterexamples are rare: the probability of encountering WL equivalent graphs in random graphs approaches 0 [2], and real-world graphs often carry rich node features that further break WL equivalences. Hence, the mere existence of theoretical counterexamples provides limited insight into a model’s practical performance on large, diverse datasets. Lastly, iso expressivity theory is graph-family agnostic. Its results are derived for the set of all possible graphs $\smash{{\mathcal{G}}^{*}}$, whereas real-world applications typically involve restricted graph families. For example, molecular datasets primarily contain planar graphs with bounded degrees. Tasks infeasible over all graphs may become solvable for standard MPNNs within these restricted graph families [13, 40]. Moreover, as shown in Tab.˜1, the basic WL test already distinguishes almost all graph pairs in popular benchmarks across different domains [54]. Therefore, iso expressivity theory cannot explain why expressivity beyond the WL test would benefit these real-world tasks or account for the performance differences between architectures.
@@ 4 Message Passing Complexity
To move beyond the limitations of expressivity theory, we propose our continuous, task-specific message passing complexity (MPC). Unlike iso expressivity theory, which only asks whether an MPNN can distinguish certain graphs in theory, MPC quantifies how difficult it is for messages to propagate through a graph to solve a given task. By accounting for the inherently lossy information propagation of real-world MPNNs, MPC can explain practical performance trends that expressivity theory misses. While many complexity measures exist in machine learning and theoretical computer science, they fail to address this unique challenge of MPNNs: propagating information effectively across graph structures. MPC specifically isolates this challenge from other general sources of difficulty in machine learning—for instance, learning high-degree polynomials is a well-understood difficulty not specific to graph learning. For clarity, we focus on node-level tasks $f_{v}$ and standard MPNNs ${\mathcal{M}}_{S}$} in this section, with generalizations to broader architectures and tasks provided in §˜B.1.
@@
Intuitively, $\alpha(x)\vDash_{X}\beta(x)$ means that $\alpha(x)$ provides sufficient information to uniquely determine $\beta(x)$. Using this, we can define a first complexity measure dependent on a specific task $f_{v}$ and individual graph G{G}. The complexity should be maximal if a task $f_{v}$ is infeasible for all $M\in{{\mathcal{M}}_{S}}$, i.e., the WL coloring ${\textup{WL}}^{L}_{v}({G})$ of $v$ in the graph G{G} contains insufficient information to deduce $f_{v}({G})$:
@@
$\textup{WLC}_{{\mathcal{M}}_{S}}(f_{v},{G})=\begin{cases}\infty&\text{if }{\textup{WL}}^{L}_{v}({G})\nvDash_{{\mathcal{G}}_{\mathcal{V}}}f_{v}({G})\\ 0&\text{else.}\end{cases}$
@@ 4.1 Weisfeiler & Leman Go Lossy
To extend our complexity measure beyond the binary characterization of the standard WL test, we need to account for the varying difficulty arising from the lossy message passing observed in practice. This difficulty inherently relies on the graph topology, i.e., the difficulty of propagating a message from one node $u$ to another node $v$ depends on 1) the number of $L$-length walks connecting $v$ and $u$ and 2) the degrees of the nodes on the walk. This can be formalized as the random walk probability ${{\bm{I}}}_{vu}^{L}$ from $v$ to $u$ with edge probabilities $\bm{I}$}} and has been connected to the amount of gradient information node $v$ receives from $u$ [47] and to the oversquashing phenomenon [19]. Correspondingly, ${{\bm{I}}}_{vu}^{L}$ serves as a measure of difficulty for the simple task of propagating a message from $u$ to $v$.
How can we adapt this measure of difficulty to arbitrary tasks? For this, we propose lossyWL, a probabilistic variant of the WL test that models the possibility of message loss. This allows us to quantify difficulty even for
Figure 2: Update step of lossyWL for node $v$. Every message $m^{l}_{a\to v}$ (green) survives independently with probability ${{\bm{I}}}_{va}$. lossyWL models the lossy message propagation observed in real-world MPNNs.
WL-distinguishable real-world graphs. Specifically, a message $m_{u\to v}^{l}$ sent from a node $u$ to its neighbor $\smash{v\in{N}_{G}(u)}$ survives with probability $\smash{{{\bm{I}}}_{vu}}$ (Fig.˜2). As a result, the color $\smash{\textup{lossyWL}_{v}^{l}}$ of a node $v$ at iteration $l$ becomes a random variable rather than a deterministic value.
@@
Let $\beta:X\to Z$ be a deterministic function and $\alpha:X\to Y$ be a probabilistic function that can be represented as a deterministic function $\alpha_{s}$ where $s$ is a seed drawn uniformly at random from a (finite) set $S$. For a fixed $x\in X$, define:
@@
MPC is based on the probability that under lossy message passing on G{G}, the output after $L$ layers contains sufficient information to deduce $f_{v}$. Intuitively, a high MPC value indicates that the task $f_{v}$ requires 1) combining information from many nodes through 2) messages of low probability, such as messages through bottlenecks. Conversely, a low MPC value means the task depends on information that remains reliably accessible even under lossy message passing conditions. Note that MPC implicitly also depends on $\mathcal{G}$}, which we omit for simplicity. While we focus here on standard MPNNs ${\mathcal{M}}_{S}$}, MPC generalizes naturally to a wide range of architectures $\mathcal{M}$} by performing lossyWL on modified message passing graphs (see §˜B.1). All in all, MPC effectively addresses the discussed limitations of iso expressivity theory, narrowing the gap between theory and practice.